A man is now 40 years old and his son is 14 years old. A number of years from now, the father will be twice as old as his son. What is this number of years?
Answer by Maths68(1474)

Man's age = 40
Sons's age = 14
Let
Number of years = x
father will be = 40 + x
Son will be = 14 + x
Twice the age of son after x years = father's age after x years
2(14+x)=40+x
28+2x=40+x
2x-x=40-28
x=12

After 12 years father will be twice as old as his son.
Check
======
Man's age = 40 +12 = 52 (which is twice of 26)
Sons's age = 14 + 12 = 26